### v1

# DDGS-CT: Direction-Disentangled Gaussian Splatting for Realistic Volume Rendering

Abstract
Digitally reconstructed radiographs (DRRs) are simulated 2D X-ray images generated from 3D CT volumes, widely used in preoperative settings but limited in intraoperative applications due to computational bottlenecks, especially for accurate but heavy physics-based Monte Carlo methods. While analytical DRR renderers offer greater efficiency, they overlook anisotropic X-ray image formation phenomena, such as Compton scattering. We present a novel approach that marries realistic physics-inspired X-ray simulation with efficient, differentiable DRR generation using 3D Gaussian splatting (3DGS). Our direction-disentangled 3DGS (DDGS) method separates the radiosity contribution into isotropic and direction-dependent components, approximating complex anisotropic interactions without intricate runtime simulations. Additionally, we adapt the 3DGS initialization to account for tomography data properties, enhancing accuracy and efficiency. Our method outperforms state-of-the-art techniques in image accuracy. Furthermore, our DDGS shows promise for intraoperative applications and inverse problems such as pose registration, delivering superior registration accuracy and runtime performance compared to analytical DRR methods.
1 Introduction
Motivation.
Digitally reconstructed radiographs (DRRs) are simulated (in silico) 2D X-ray images rendered from 3D computational tomography (CT) volumes. While DRRs are widely utilized in preoperative settings, such as optimizing dose delivery and in radiation oncology Lohr et al. (1997); Yang et al. (2000); Bollet et al. (2003), their potential intraoperative applications remain underexplored due to computational bottlenecks and naive modeling capability Adler Jr et al. (1999); Peters (2001). For instance, real-time multimodal registration for image-guided procedures is currently impractical because of the time-consuming process of generating DRRs and integrating them into slice-to-volume registration Van Der Bom et al. (2011).
The advent of GPU-accelerated computing has significantly improved the efficiency of radiography simulators Badal and Badano (2009); Vidal et al. (2009); Biguri et al. (2016), though they still fall short of meeting the stringent requirements of real-time applications. Recent developments in inverse graphics have led to the creation of differentiable DRR renderers (e.g., DiffDRR Gopalakrishnan and Golland (2022), X-Gaussian Cai et al. (2024), GaSpCT Nikolakakis et al. (2024)), which show promise for inverse problems such as 2D/3D CT image registration Gopalakrishnan et al. (2023). However, these methods do not fully capture the complexities of X-ray image formation. They typically use computationally efficient ray-tracing techniques that model the attenuated photon fluence at each detector pixel by accumulating Hounsfield unit (HU) values along the 3D view-ray through the voxelized CT volume. This approach, while fast, cannot account for complex noise-inducing physical effects such as Compton scattering and beam hardening Hasegawa (1990), crucial for accurate X-ray imaging.
Physics-based Monte-Carlo simulations Agostinelli et al. (2003); Jan et al. (2004); Badal and Badano (2009); Bert et al. (2013) do not suffer from these limitations. Relying on the decomposition of CT volumes into realistic materials via HU thresholding, they accurately simulate single-photon transport and probabilistically evaluate photon-matter interactions (photoelectric effect, Compton scattering, Rayleigh scattering, etc.) to determine the final attenuation and hit-region of each photon. This approach provides a more accurate representation of X-ray image formation; but the computational intensity of these methods—requiring >$10^{8}$ iterations to render an image—makes them impractical for real-time applications and inverse problems.
In this paper, we propose a novel, more elegant balance between (a) realistic physics-inspired X-ray simulation and (b) efficient, differentiable DRR generation, as illustrated in Figure  1.
Contributions.
Our approach leverages 3D Gaussian splatting (3DGS) Kerbl et al. (2023) for efficient DRR generation from CT volumes, similar to X-Gaussian Cai et al. (2024) and GaSpCT Nikolakakis et al. (2024). However, we introduce a crucial adaptation: our direction-disentangled 3DGS (DDGS) method accounts for the dependencies of X-ray image formation on light directionality. Specifically, we decompose the radiosity contribution of 3D regions into two independent components: a non-directional component modeling the isotropic X-ray interactions with the corresponding material, and a direction-dependent component approximating higher-dimensional anisotropic interactions, such as scattering. Unlike the more generic view-dependent spherical-harmonics decomposition adopted by common 3DGS solutions Fridovich-Keil et al. (2022), ours is tuned to the specific behaviors of high-energy photons. This formulation allows our method to implicitly learn and reproduce the effects of scattering on X-ray images without the need for complex photon-transport simulations at runtime.
Additionally, we adapt the initialization phase of 3DGS to account for the properties of tomography data. We adopt a geometry- and intensity-based dual strategy to sample 3D points that correspond to: (a) contact regions between different anatomical elements, where photon transport is likely to show discontinuities; and (b) homogeneous regions, where the primary attenuation contribution can be sparsely modeled but could benefit from view-dependent noise modeling.
We demonstrate that our solution achieves higher image accuracy compared to state-of-the-art methods across various medical benchmarks Additionally, we showcase the efficient application of our method to inverse problems, such as pose registration, highlighting its potential for real-time intraoperative use cases. Our DDGS achieves better registration accuracy and runtime performance than analytical DRR methods.
Figure 1: Proposed pipeline of Direction-Disentangled 3D Gaussian Splatting (DDGS).
2 Related Work
The generation of DRRs and the simulation of X-ray images have been subjects of extensive research due to their critical roles in medical imaging, particularly in radiotherapy planning and diagnostic imaging Lohr et al. (1997); Yang et al. (2000); Bollet et al. (2003); Zhang et al. (2022). In this section, we provide some insight on X-ray imaging, traditional simulation tools, and the recent integration of machine-learning techniques.
X-ray Image Simulation.
X-ray imaging relies on the interactions of high-energy photons—emitted by the X-ray source of the scanner—with the matter (e.g., body part) placed in between the X-ray source of the scanner and its radiation-sensitive detector cells (dexels). Imaging contrast arises from differential absorption and scattering of X-ray photons by tissues, primarily through the photoelectric effect, Compton scattering, and, to a lesser extent, beam hardening for polychromatic scanners. The photoelectric effect involves the absorption of an X-ray photon by an atom and is more pronounced in higher atomic number materials like bone, causing them to appear whiter on X-ray images [^1] X-ray film is historically white, turning black when exposed to high-energy photon.; whereas Compton scattering occurs when an X-ray photon collides with an outer electron, resulting in the photon being deflected and losing energy, which contributes to image noise and reduced contrast Hasegawa (1990); Seibert and Boone (2005).
A variety of Monte Carlo (MC) simulation suites Agostinelli et al. (2003); Jan et al. (2004); Badal and Badano (2009); Bert et al. (2013) model these X-ray interactions and light transport. They trace the paths of a large number of individual photons as they undergo probabilistic interactions within the material. By simulating numerous photon trajectories, Monte Carlo methods account for the random nature of photoelectric absorption and Compton scattering, as well as secondary effects such as fluorescence and electron transport. These MC solutions produce highly realistic simulations of radiographic images, though at a high computational cost which makes them inadequate for most DRR use-cases. Furthermore, for these tools to render realistic DRRs from CT volumes, one must first pre-process the CT data to assign proper materials and their corresponding physical properties to each voxel, which is essential for realistic simulation. This step is often approximated by decomposing the volume into a small set of predefined materials according to user-provided HU ranges (e.g., “air” material assigned to voxels with HU value <800<-800< - 800, “lung” assigned to voxels in range ]800,200]]-800,-200]] - 800 , - 200 ], “fat” for ]200,100]]-200,-100]] - 200 , - 100 ], etc.). This process can be time-consuming and prone to errors, affecting the fidelity of MC-based DRR generation Dorgham et al. (2020).
Efficient and Differentiable DRR Generation.
Analytical DRR rendering methods Vidal et al. (2009); Vidal and Villard (2016); Biguri et al. (2016) have been developed in parallel to the aforementioned MC suites for real-time or near-real-time applications. These methods have a much lower computational footprint, as they do not simulate individual photon interactions but instead use mathematical models to directly calculate the primary paths and approximated attenuations of X-rays. The seminal work of Siddon (1985) on computing line integrals through a discretized CT volume is still at the core of most recent DRR solutions, e.g., which proposed differentiable formulation of the rendering steps Gopalakrishnan and Golland (2022) or integrated neural networks to add realistic noise to the output Unberath et al. (2018). However, the inherent approximations in these methods can affect their accuracy, particularly in modeling artifacts present in heterogeneous tissues or complex anatomical structures, e.g., caused by scattering (non-primary light contributions).
Recently, researchers Cai et al. (2024); Nikolakakis et al. (2024) have tried to apply 3D Gaussian splatting (3DGS) techniques Kerbl et al. (2023) to DRR rendering, i.e., optimizing a cloud of 3D Gaussians to approximate voxel data, thereby enabling faster rendering with minimal accuracy loss. However, these early solutions overlooked the specific nuances of X-ray imaging, directly applying 3DGS methods designed for natural imaging Nikolakakis et al. (2024) or oversimplifying the physical properties of X-ray attenuation Nikolakakis et al. (2024), leading to suboptimal modeling.
In this paper, we propose a novel formulation of 3D Gaussian splatting, tuned to more efficiently approximate X-ray imaging (time- and quality-wise).
3 Methodology
3.1 Preliminaries
Gaussian Splatting.
3DGS Kerbl et al. (2023) is a point-based rendering technique using 3D Gaussians to explicitly represent scenes in a more compact manner than volumetric representations. Each Gaussian $g_{i}=\{\bm{\mu}_{i},\bm{\Sigma}_{i},\bm{\alpha}_{i},\bm{f}_{i}\}$ is defined by its mean position $\bm{\mu}_{i}\in\mathbb{R}^{3}$, covariance matrix $\Sigma_{i}\in\mathbb{R}^{3\times 3}$ (usually decomposed into separate scale and rotation parameters), opacity $\bm{\alpha}_{i}\in\mathbb{R}$, and radiance properties $\bm{f}_{i}\in\mathbb{R}^{k}$, where $k$ depends on the light contribution model adopted. E.g., $k=(L+1)^{2}\times 3$ for anisotropic, view-dependent RGB radiance decomposed into spherical harmonics (SH) of degree $L$, a model commonly used for natural-imaging applications. Images are rendered by casting view-rays through each pixel $p$ into the scene and alpha-blending the Gaussian contributions to the final ray color $C$, as:
$C(p)=\sum_{j\in N}{c_{j}\sigma_{j}\prod_{l=1}^{j-1}{(1-\sigma_{l})}}\;\text{ with }\sigma_{i}=\bm{\alpha}_{i}e^{-\frac{1}{2}(p-\widehat{\bm{\mu}}_{i})\widehat{\bm{\Sigma}}_{i}^{-1}(p-\widehat{\bm{\mu}}_{i})},$ (1)
where $N$ is the number of points sampled along the view-ray, $c_{j}$ is the radiance (computed from $\bm{f}_{j}$, e.g., via SH) of the $j$th Gaussian on the ray, and $\widehat{\bm{\mu}}_{j}$ and $\widehat{\bm{\Sigma}}_{j}$ are the image-plane projections of $\bm{\mu}_{j}$ and $\bm{\Sigma}_{j}$.
Since the aforementioned rasterization process is differentiable w.r.t. to the Gaussian parameters, the 3DGS representation of a scene can be learned via gradient-descent, given a set of 2D scene observations $\bm{I}_{\text{GT}}$ and their corresponding camera parameters. A combination of $\ell_{1}$ and SSIM Wang et al. (2004) loss functions are commonly-adopted as criterion for this iterative optimization process Kerbl et al. (2023):
$\mathcal{L}=(1-\lambda)\ell_{1}{(\bm{I}_{\text{GT}},\bm{I})}+\lambda\operatorname{SSIM}{(\bm{I}_{\text{GT}},\bm{I})},$ (2)
with $\bm{I}$ the rendered images (based on provided camera parameters), and $\lambda$ a loss-weighting hyper-parameter. The initial state is typically obtained by sampling relevant 3D points in the scene domain for the Gaussians, e.g., from a structured-from-motion (SfM) point-cloud Ullman (1979); Schonberger and Frahm (2016).
Application to DRR.
The volumetric rendering performed by 3DGS methods is conceptually similar to the one performed by analytical DRR methods Siddon (1985); Gopalakrishnan and Golland (2022), i.e., aggregating the attenuation values of light-rays cast through the voxelized CT volume. This similarity has recently motivated researchers Cai et al. (2024); Nikolakakis et al. (2024) to apply 3DGS models to DRR applications, driven by the compactness of 3DGS representations compared to voxel data (addressing the $\mathcal{O}(n^{3})$ complexity inherent to voxel rendering). By optimizing a 3DGS model to approximate the visual properties of the CT volume, the resulting representation can render accurate DRRs much faster than traditional methods, making it more suitable for real-time intra-operative visualization. Consequently, the use of traditional, slower rendering solutions can be limited to generating DRR targets ($\bm{I}_{\text{GT}}$) for the offline (i.e., pre-operative) training.
However, these early 3DGS-for-DRR solutions Siddon (1985); Gopalakrishnan and Golland (2022) do not account for noise-inducing photon interactions (e.g., scattering) when applying the analytical methods Sharp et al. (2010); Biguri et al. (2016) to render their training data. Since scatter-free DRRs are not affected by ray directions (i.e., the attenuation at each point is naively considered isotropic), Cai et al. (2024) simplified the radio-intensity function of each Gaussian to $c_{i}=\operatorname{sigmoid}{(\bm{b}\cdot\bm{f}_{i})}$, where $\bm{b}\in\mathbb{R}^{k}$ is a direction-independent optimizable basis vector shared by all splats. Furthermore, their Gaussian point-cloud is initialized via evenly-spaced Cai et al. (2024) or random-uniform Nikolakakis et al. (2024) sampling of the scanned 3D space, i.e., ignoring relevant geometrical and material properties of the target CT data. Inadequate initialization strategies negatively impact the convergence of 3DGS models and the resulting image accuracy Jung et al. (2024). In this work, we propose to reformulate both the initialization strategy and the radiance function of the 3D Gaussians to address these limitations.
3.2 Disentanglement of Isotropic and Anisotropic 3D Gaussians
We argue that, while X-ray scattering models are incompatible with the 3DGS rendering formulation (secondary rays would have to be cast, increasing the computational footprint exponentially), their anisotropic impact on X-ray imaging can be approximated to some extent by direction-dependent radiance functions. However, modeling this high-dimensional residual contribution is only meaningful if it appears in the training DRRs, i.e., if the target radiographs are rendered using physics-inspired simulation tools. This is not always possible, e.g., if the mapping from HU values to materials for the target CT scans is not provided. Ideally, in such cases, when only scatter-free DRRs can be generated, the radiance function should gracefully degrade to a lighter anisotropic formulation (e.g., similar to Cai et al. (2024)), to avoid unnecessary computations that could impact the model optimization, as well as its runtime performance. Moreover, X-ray interactions with matter themselves involve both isotropic (e.g., photoelectric absorption/fluorescence, Rayleigh scattering) and anisotropic (e.g., Compton scattering) phenomena Hasegawa (1990); Migliori . Therefore, for the sake of both modularity and accuracy, we propose to decompose our DRR representation into isotropic Gaussians $g_{i}^{\text{iso}}$ and anisotropic, direction-dependent ones $g_{j}^{\text{dir}}$. We formulate their respective radiosity functions as:
$c_{i}^{\text{iso}}=\operatorname{sigmoid}{(\bm{b}^{\text{iso}}\cdot\bm{f}_{i}^{\text{iso}})}\;\;\;\;;\;\;\;\;c_{j}^{\text{dir}}=\operatorname{sigmoid}{\left(Y_{1..L}(\theta,\phi)\cdot\bm{B}^{\text{dir}}\bm{f}_{j}^{\text{dir}}\right)},$ (3)
where $\bm{f}_{i}^{\text{iso}},\bm{f}_{j}^{\text{dir}}\in\mathbb{R}^{k}$ are feature vectors respectively encoding the contribution of isotropic and anisotropic Gaussians, $\bm{b}^{\text{iso}}\in\mathbb{R}^{k}$ is a global isotropic basis vector (similar to Cai et al. (2024)), and $\bm{B}^{\text{dir}}\in\mathbb{R}^{k_{L}\times k}$ is a second, anisotropic basis matrix that interacts with the $L$-degree spherical-decomposition $Y:\mathbb{R}^{2}\mapsto\mathbb{R}^{k_{L}}$ applied to the ray angles $\theta,\phi$ without constant term (degree $=$ 0), resulting in $k_{L}=L(L+2)$.
It is important to note that, while the usual spherical-harmonics representation for Gaussian splatting also contains isotropic (degree $=$ 0) and anisotropic (degree $>$ 0) components, the model assumes that both contributions are co-located (i.e., they belong to the same 3D Gaussian $g_{i}$ of position $\bm{\mu}_{i}$ and covariance $\bm{\Sigma}_{i}$). To account for the complexity of X-ray light transport, we relax this co-location constraint, i.e., modeling isotropic and anisotropic contributions via distinct 3D Gaussians ($g_{i}^{\text{iso}}$ and $g_{j}^{\text{dir}}$). We demonstrate in our evaluation that this disentanglement results in both higher image quality and lighter representation.
3.3 Initialization via Radiodensity-Aware Dual Sampling (RADS)
Our second main contribution targets the initialization of the point-cloud that supports the Gaussian-mixture modeling of the 3D data. Since 3DGS was defined for natural imaging (e.g., ignoring sub-surface light transport), common initialization strategies limit their sampling to scene-surface points (SfM- Kerbl et al. (2023) or depth-guided Niemeyer et al. (2024) subsampling). Such sampling of the 3D space is, however, inadequate for volumetric data, ignoring possibly salient regions. While uniform sampling of the CT space could be considered Cai et al. (2024); Nikolakakis et al. (2024), we argue that such a strategy is also suboptimal for anatomical data, discarding domain-relevant properties. We demonstrate empirically (c.f. Section 4) that uniform sampling results in slower convergence of the 3DGS representation and degraded runtime performance.
Therefore, we propose a novel twofold sampling strategy that accounts for the specific nature of CT data and for our dual isotropic/directional 3DGS model. Based on the radiodensity values contained in the target CT volumes, our solution samples two set of 3D points $\mathbf{P}^{\text{mc}}\in\mathbb{R}^{n_{1}\times 3}$ and $\mathbf{P}^{\text{dw}}\in\mathbb{R}^{n_{2}\times 3}$ with distinct, complementary distributions ($n_{1},n_{2}$ scalar hyper-parameters).
Not unlike Niemeyer et al. (2024), the first set of points is sampled by applying the marching-cubes algorithm Lorensen and Cline (1998) to extract points at the interface between materials with different physical properties (i.e., distinct HU responses in the CT volume). These interface regions typically need careful modeling when simulating X-ray imaging, as they result in discontinuous photon transport.
To complement this first set of points, the second set is semi-randomly sampled from the voxel centroids, according to a uniform distribution weighted by the voxels’ radiodensities. I.e., voxels with higher radiodensity, contributing to the X-ray attenuation, are more likely to be picked to initialize the Gaussians. In other words, this radiodensity-weighted strategy identifies additional points within the homogeneous regions of the CT volume that also contributes to the imaging process.
Based on their distinct properties, we assign the entire $\mathbf{P}^{\text{mc}}$ to initialize the isotropic Gaussians, whereas we equally split $\mathbf{P}^{\text{dw}}$ to initialize both isotropic and direction-dependent Gaussians. All in all, our dual radiodensity-aware strategy improves the initialization of the 3DGS representation and is more explainable than prior work (see Figure 2).
Figure 2: Illustration of the different sampling strategies for 3DGS initialization.
4 Experiments
4.1 Experimental Protocol
Implementation.
For evaluation, we set the decomposition degree to $L$=1, the feature dimension to $k$=8, and initial cloud sizes to $n_{1}$=15,000 and $n_{2}$=10,000. We apply a loss weight $\lambda=0.2$. We choose the default threshold level (i.e., the average of the minimum and maximum volume) for the marching-cubes algorithm. DDGS training is performed on a single NVIDIA A100 using an Adam optimizer Kingma and Ba (2014) with a learning rate of $1.25\times 10^{-4}$ for $\bm{b}^{\text{iso}}$ and $\bm{B}^{\text{dir}}$ and $2.5\times 10^{-3}$ for $\bm{f}^{\text{iso}}$ and $\bm{f}^{\text{dir}}$. Default 3DGS Kerbl et al. (2023) learning rates are applied to the remaining parameters.
Datasets.
We consider 4 datasets. (1) NAF-CT Zha et al. (2022) includes four CT images of abdomen, chest, foot, and jaw. For each, we adopt TIGRE Biguri et al. (2016) to sample 50 evenly-distributed projections for training and 50 randomly-distributed ones for testing, in the range of [90°- 90  °, 90°90  °] (scatter-free DRRs). (2) CTPelvic1K Liu et al. (2021) is a large dataset of pelvic CT images. We consider the first 10 scans of the sub-dataset6 in our experiments. We adopt DeepDRR Unberath et al. (2018) to generate 60 training (evenly-distributed) and 60 (randomly-distributed) testing DRRs (in the range of [60°- 60  °, 60°60  °]), taking advantage of the scattering-modeling capability of DeepDRR to render realistic X-ray data. (3) Ljubljana Pernus et al. (2013) is a clinical dataset of 10 patients undergoing neurovascular surgery. Each patient underwent one CT and two X-ray angiography scans. The pose of each X-ray is also provided. Following DiffPose strategy Gopalakrishnan et al. (2023), we randomly sample 900 projections for training and 100 for testing. (4) Finally, we provide some additional qualitative results on DeepFluoro Grupp et al. (2020), a collection of pelvic X-rays and CT images from six cadavers.
Metrics.
We evaluate our method in two settings: novel-view synthesis and intraoperative 2D/3D image registration (i.e., pose estimation). For novel-view synthesis, we use the standard PSNR and SSIM as metrics. For intraoperative 2D/3D image registration, we consider the error both in terms of rotation (angular distance) and translation (Euclidean distance). We further measure the clinically-relevant registration accuracy of key anatomical landmarks for Ljubljana scans Pernus et al. (2013) where they are provided. Following Gopalakrishnan et al. (2023), we compute the target registration error (TRE) w.r.t. the positioning of these landmarks after registration.
4.2 Novel-View Synthesis
We first validate our contributions in terms of the quality of rendered DRRs and compactness of our dual Gaussian-mixture representation.
Table 1: Comparison of Gaussian splatting-based DRR rendering techniques, on 2 datasets.
3DGS ($L=1$) X-Gaussian ($k=32$) DDGS (Ours, $L=1,k=8$)
[…]
NAF-CT
abdomen 11,149 47.43 0.994 10,802 47.17 0.993 13,928 48.09 0.994
chest 13,669 44.42 0.988 16,568 43.33 0.987 13,533 44.50 0.989
foot 8,616 44.51 0.984 10,909 44.34 0.985 10,786 44.70 0.985
jaw 17,902 40.47 0.973 22,318 40.02 0.972 19,665 40.57 0.974
avg 12,834 44.21 0.985 15,149 43.72 0.984 14,478 44.47 0.986
CTPelvic1K-Dataset6
001 53,988 35.40 0.971 50,059 36.81 0.979 41,947 37.88 0.984
002 49,099 37.03 0.982 48,113 37.79 0.986 43,933 38.43 0.988
003 60,755 35.73 0.973 59,822 36.46 0.977 48,536 38.28 0.983
004 42,349 38.87 0.982 37,243 39.84 0.985 39,176 40.35 0.986
005 42,482 39.37 0.984 46,014 39.30 0.984 39,570 40.36 0.987
006 53,832 37.14 0.983 57,197 37.42 0.983 47,398 38.26 0.986
007 45,360 37.09 0.980 48,454 37.62 0.983 41,032 38.87 0.987
008 51,211 38.03 0.980 45,750 38.70 0.982 43,577 39.81 0.986
009 44,060 38.19 0.981 41,279 38.92 0.985 41,389 39.87 0.987
010 38,691 37.80 0.984 40,030 37.55 0.985 39,691 38.73 0.987
avg 48,183 37.47 0.980 47,396 38.04 0.983 42,625 39.08 0.986
Comparative Evaluation.
We compare to 3DGS Kerbl et al. (2023) and X-Gaussian Cai et al. (2024) on the NAF-CT Zha et al. (2022) and CTPelvic1K Liu et al. (2021) datasets for the novel-view synthesis. For each method, we adopt the hyper-parameters recommended by their respective authors (e.g., $L=1$ and 333 for normal 3DGS Kerbl et al. (2023) and $k=8$ and 323232 for X-Gaussian Cai et al. (2024)).
Table 2: Novel-view synthesis evaluation on CTPelvic1K data.
500 2000 7000 15000 30000
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
DDGS (Ours)
($L=1,k=8$)
001 25.77 0.908 30.46 0.944 35.18 0.975 36.91 0.981 38.04 0.984
002 27.64 0.921 31.60 0.954 35.33 0.977 36.90 0.984 38.30 0.987
DDGS (Ours)
($L=3,k=8$)
001 26.20 0.909 31.50 0.950 35.65 0.978 37.68 0.983 38.15 0.985
002 27.87 0.923 32.23 0.956 36.02 0.979 37.87 0.986 38.97 0.989
3DGS
($L=1$)
001 23.56 0.873 27.52 0.914 31.66 0.951 33.64 0.963 35.36 0.970
002 24.55 0.891 29.17 0.932 32.82 0.965 35.33 0.977 37.25 0.982
3DGS
($L=3$)
001 23.55 0.872 27.54 0.914 33.34 0.962 36.01 0.975 37.12 0.980
002 24.55 0.891 28.96 0.931 34.50 0.973 37.72 0.985 38.75 0.988
X-Gaussian
($k=8$)
001 17.31 0.794 24.79 0.877 29.00 0.923 30.72 0.933 33.39 0.947
002 13.45 0.742 27.19 0.904 32.72 0.962 34.52 0.973 36.04 0.978
X-Gaussian
($k=32$)
001 26.34 0.854 31.28 0.950 33.56 0.968 35.07 0.974 36.83 0.979
002 28.32 0.868 33.44 0.966 35.85 0.980 36.55 0.984 37.89 0.986
As shown in Table 1, for NAF-CT Zha et al. (2022) (without scatter-related noise), our DDGS outperforms both 3DGS and X-Gaussian in terms of PSNR and SSIM. For CTPelvic1K Liu et al. (2021) (with scatter), our DDGS significantly outperforms 3DGS and X-Gaussian in both PSNR and SSIM while using considerably fewer points. Therefore, our DDGS method is more effective in simulating realistic X-ray images from CT scans.
Furthermore, we compare our DDGS (with degrees $L=1$ and $L=3$ and feature dimension $k=8$) with traditional 3DGS (degrees $L=1$ and $L=3$) and X-Gaussian (feature dimensions $k=8$ and $k=32$) on the 001 and 002 data of CTPelvic1K Liu et al. (2021) at iterations of 500, 2000, 7000, 15,000, and 30,000, as shown in Table 2. Our model with $L=1$ and $k=8$ performs comparably to 3DGS with $L=3$ and outperforms X-Gaussian with $k=32$ after 15,000 iterations. Our model with $L=3$ achieves the best performance overall. Figure 3 presents a qualitative comparison of the testing set, illustrating both the synthetic views and the generated 3D Gaussian points.
Figure 3: Visualization of the Gaussian cloud optimization for different methods.
Table 3: Ablation study w.r.t. the proposed disentangled isotropic/anisotropic representations.
DDGS (Ours) direct-entangled direct-independent direct-dependent 3DGS-disentangled
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
001 38.04 0.984 36.17 0.977 35.93 0.976 32.48 0.942 35.32 0.971
002 38.30 0.987 37.32 0.984 36.96 0.983 33.04 0.953 35.38 0.973
Table 4: Impact of initialization strategies on final image quality, measured CTPelvic1K scans.
001 002 003 004 005 006 007 008 009 010 avg
Radiodensity-based (Ours) 37.88 38.43 38.28 40.35 40.36 38.26 38.87 39.81 39.87 38.73 39.08
Random (uniform dis.) 37.17 38.36 38.15 40.09 39.97 38.36 38.69 39.37 39.85 38.84 38.89
Even sampling 37.33 38.51 37.83 39.58 39.12 38.27 38.87 39.88 39.46 38.95 38.78
Table 5: Impact of $k$ dimensionality on image quality, measured on 2 CTPelvic1K scans (001, 002).
$k=1$ $k=4$ $k=8$ $k=16$ $k=32$
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
001 34.60 0.965 36.81 0.979 38.04 0.984 38.41 0.986 38.27 0.987
002 33.66 0.964 37.08 0.982 38.30 0.987 38.76 0.989 38.89 0.989
Figure 4: Visualization of the isotropic and anisotropic X-ray contributions to the rendered DRRs.
Figure 5: Illustration of the pose-registration convergence using various differentiable DRR renderers.
Table 6: Accuracy of image registration on the CTPelvic1K Liu et al. (2021) dataset.
3DGS X-Gaussian ($k$=8) X-Gaussian ($k$=32) DDGS (Ours)
001 002 003 001 002 003 001 002 003 001 002 003
Rot. error (deg) 0.134 0.062 0.229 0.326 0.163 0.312 0.097 0.070 0.266 0.067 0.067 0.152
Trans. error (mm) 1.788 1.163 3.108 4.037 2.144 3.781 1.378 1.288 3.483 1.285 1.092 2.180
Table 7: Registration evaluation w.r.t. anatomical landmark accuracy and time efficiency, on Ljubljana.
Method DRR render time (ms) ↓ total optimization time (s) ↓ TRE (mm) ↓
mean med std mean med std mean med std
DiffDRR 30.88 31.51 3.25 431.98 316.54 261.70 2.14 0.72 5.40
DDGS 7.46 6.94 2.67 121.45 39.93 117.67 0.80 0.65 0.51
Ablation Study.
Table 3 demonstrates the effectiveness of our direction-disentangled representation for 3D Gaussians. For several settings, we observe performance degradation when learning the direction-dependent and direction-independent components in the same 3D Gaussians (direct-entangled), learning the direction-independent 3D Gaussians alone (direct-independent), or learning the direction-dependent 3D Gaussians alone (direct-dependent). Additionally, we decouple the direction-dependent and direction-independent components in 3DGS Kerbl et al. (2023) (3DGS-disentangled), which also showed lower performance compared to our learnable features. Figure 4 shows the rendered views from disentangled 3D Gaussians, highlighting that the direction-dependent 3D Gaussians focus more on bones or anatomical structures, while the direction-independent 3D Gaussians focus more on the background.
Table 4 shows the effectiveness of our proposed initialization scheme (radiodensity-aware) compared to other initialization methods. Our scheme outperforms both the uniformly random sampling used in 3DGS Kerbl et al. (2023) and the even sampling method used in X-Gaussian Cai et al. (2024).
At last, we conduct experiments on the impact of feature dimension $k$, as shown in Table 5. Our results indicate that increasing the feature dimension generally improves overall performance. However, the improvement becomes marginal once the feature dimension exceeds $k=16$.
4.3 Application to Downstream Task (2D/3D CT Image Registration)
We validate our solution in terms of applicability to downstream tasks, considering here the registration of intraoperative 2D X-ray images to preoperative 3D CT volumes; a task crucial to a variety of clinical applications.
We choose the first five testing images on the 001, 002, and 003 CT data of CTPelvic1K Liu et al. (2021) for image registration. Considering this task as an optimization-based problem that can be solved via inverse graphics (i.e., rendering synthetic images based on predicted poses and comparing to the target image, then backpropagating the difference to improve the pose prediction), we adopt the framework from iComMA Sun et al. (2023), replacing their image rendered by DDGS. We use Adam optimizer with a learning rate of 0.05. We use the isocenter pose as the initialization pose. As shown in Table 6 and Figure 5, our DDGS achieves the lowest rotation and translation errors.
Finally, we also consider the experimental protocol proposed in DiffPose Gopalakrishnan et al. (2023), where the authors rely on a pretrained pose-regression CNN to get a first rough estimation of the scanner pose, before refining said pose via a gradient-descent based optimization, leveraging their differentiable DRR renderer Gopalakrishnan and Golland (2022). We adopt their pose-initialization network and testbed for the Ljubljana dataset, replacing their DiffDRR analytical renderer Gopalakrishnan and Golland (2022) by our DDGS, and compare to them in terms of runtime performance and landmark-registration accuracy. The results, reported in Table 7, highlight the convergence boost brought by our lighter solution.
5 Conclusion
We proposed DDGS, a novel 3DGS-based representation for efficient and realistic DRR generation from CT volumes. We extended traditional 3DGS by considering light directionality, crucial for accurate X-ray image modeling. By decomposing radiosity into non-directional and direction-dependent components, DDGS captures isotropic interactions and anisotropic effects. Unlike generic view-dependent spherical-harmonics decomposition, DDGS implicitly learns scattering effects, substantially improving the speed and accuracy of downstream clinical applications such as intraoperative 2D/3D registration.
Limitations. It should be noted that, similar to previous methods, our model expects an offline DRR renderer to produce realistic ground truth to train DDGS, which can be a limitation for some scenarios (e.g., when such renderer or certain parameters required for realistic rendering are not available). Our direction-dependent function relating to anisotropic X-ray effects may also fail to approximate complex multi-bounce light scenarios. Recent developments in $N$-dimensional Gaussian splatting Diolatzis et al. (2024) show promising potential for modeling such cases.
Societal Impact. The above limitation may restrict the adoption of DDGS to specific use cases, and our method should undergo a clinical evaluation of the generated images in terms of anatomical accuracy. Nonetheless, we believe that our proposed method can positively impact both the computer vision and medical communities by providing a more efficient and versatile DRR tool.